OpenStudy (anonymous):

Which ratio is not equivalent to 45/120? A.3/8 B.5/12 C.9/24 D.90/240

OpenStudy (anonymous):

@sbta2

OpenStudy (anonymous):

well :) 5x9=45 but 5x12=60...so 5/12 does not equal 45/120

OpenStudy (texaschic101):

try reducing the fraction 45/120 45/120 = 9/24 = 3/8 9/24 also = 90/240 so I agree with sbta.....5/12 does not belong

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

thats what we are here for

OpenStudy (anonymous):

Which ratio is proportional to Proportion 3/5 ? A.9/10 B.12/20 C.15/20 D.5/7

OpenStudy (anonymous):

i think its a

OpenStudy (anonymous):

Can you say why you think it's a? You need to find one number that when multiplied to 3 and to 5 that gives you one of the answers.

OpenStudy (texaschic101):

3/5 x 4/4 = 12/20